### v1

Abstract    Smoluchowski-Kramers approximation in rough path topology with state-dependent damping is explored. The second-order Langevin equation has a form of fast-slow system after suitable change-of-variable, and then its solution is lifted as a rough path in a natural manner. Moment estimates of both the original path and the lift are given, followed by which, averaging technique and convergence theorem in rough path topology are used to pass the limit.
Keywords    Smoluchowski–Kramers approximation, Averaging, Singular pertubation, Rough path theory
2020 Mathematics Subject Classification   60L90, 60H10
1 Introduction
Consider the state-dependent second-order Langevin equation
$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M(X^{\epsilon}_{t})\dot{X}^{\epsilon}_{t}=F(X^{\epsilon}_{t})+\dot{W}_{t},\kern 5.0ptX^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0,$ (1.1)
defined on a complete stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},P),$ where $0<T<\infty$ is fixed, $M$ is a matrix-valued function and $F$ is a non-linearity, detailed in Section 2. We are interested in the limit of $X^{\epsilon}$ as $\epsilon\to 0,$ which is called Smoluchowski–Kramers (SK) approximation. There is plenty of research for the case $M$ is a constant [2, 3, 33, 16, 30, 7], and in this case, it is not surprising that $X^{\epsilon}\to X$ in some sense with
$M\dot{X}_{t}=F(X_{t})+\dot{W}_{t},\kern 5.0ptX_{0}=0,$
which can be obtained formally by taking $\epsilon=0$ in (1.1). However, when $M$ is not a constant, that is, it is state-dependent, the limit cannot be obtained by simply dropping the small mass term. In fact, under certain assumptions, Hottovy et al.[12] showed that
$\lim_{\epsilon\rightarrow 0}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}-X_{t}|^{2}=0,$
where
$dX_{t}=S(X_{t})dt+M^{-1}(X_{t})F(X_{t})dt+M^{-1}(X_{t})dW_{t},\kern 5.0ptX_{0}=0,$ (1.2)
and
$S_{j}(x)=\sum_{k,l}(\partial_{x_{l}}M^{-1})_{jk}(x)J_{kl}(x),$
with $J$ the solution of Lyapunov equation
$M(x)J(x)+J(x)M^{\top}(x)=\mathrm{id}.$ (1.3)
Since then, there are fruitful work on SK approximation with state-dependent damping. For example, Wang et al.[27] explore the limit by averaging on the Fokker-Planck equation associated with the Langevin equation. Cerrai and Xi[4] consider the infinite dimensional case. For more research in SK approximation in state dependent case, we refer to [5, 11, 17, 28].
SK approximation is naturally linked to stochastic averaging, initiated by Khaminskii [13]. Numerous work concerning about stochastic averaging appeared in recent years, see e.g. [14, 29, 32, 31]. There is also research about averaging principle in rough path framework, see for example in [15, 25]. In fact, by suitable change-of-variable, (1.1) is transformed to (2.3), which has a form of fast-slow system [6], making it possible to find its limit by averaging method. A Poisson equation technique[21, 22, 23] is employed to show the convergence.
Rough path theory is initially proposed by Lyons[19]. Although a deterministic theory, it is proved a powerful tool in stochastic analysis. See monographs [18, 20, 9, 8] for systematic display of rough path theory. In this paper, we aim to find the small mass limit in rough path topology. Friz et al [7] identify the limit in rough path topology in the case that $M$ is a constant matrix and $F\equiv 0$. They found that, the limit of the "level-2" rough path, converges to a non-trivial limit. In particular, when the dimension is $1$, the Itô lift of the original path, converges to the Stratonovich lift of the limit path, which is reminiscent of Wong–Zakai approximation [26]. They also gave a conjecture for the case when $M$ is not constant, without detailed proof. In this paper, the system with both state-dependent friction and non-linearity is explored. We identify the limit in the rough path topology, and give rigorous proof. This also generalizes the result in [12], in the sense that in our main theorem, the "level-1" path converges in $p$-th moment of $\alpha$-Hölder norm for all $1\leq p<\infty$.
The rest of this paper is organized as follows. In Section 2, we recall some basic rough path theory, list some notations, and present our main result. In Section 3, we give moment estimate in rough path norm. In Section 4, we display an averaging principle for specific fast-slow system, which is used to pass the limit. After these preparation, we prove the main theorem in Section 5.
2 Preliminary and Main Result
We recall some basic definitions of rough path theory [8, e.g.]. Let $X:[0,T]\to\mathbb{R}^{d},$ $\mathbb{X}:\Delta_{[0,T]}\to\mathbb{R}^{d\times d}$, where
$\Delta_{[0,T]}:=\{(s,t)\in[0,T]^{2}:s<t\}.$
Given $\frac{1}{3}<\alpha\leq\frac{1}{2},$ a map $\mathbf{X}=(X,\mathbb{X})$ is called an $\alpha$-Hölder rough path over $\mathbb{R}^{d}$, if the following conditions hold:
(i) $X||_{\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|X_{t}-X_{s}|}{|t-s|^{\alpha}}<\infty$
(ii) $\mathbb{X}||_{2\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|\mathbb{X}_{s,t}|}{|t-s|^{2\alpha}}<\infty$
(iii) (Chen relation) $\mathbb{X}_{s,u}=\mathbb{X}_{s,t}+\mathbb{X}_{t,u}+X_{s,t}\otimes X_{t,u},\kern 5.0pt\forall 0\leq s\leq t\leq u\leq T.$
The set of all $\alpha$-Hölder rough path over $\mathbb{R}^{d}$ is denoted as $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}).$ Given $\mathbf{X},\mathbf{Y}\in\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}),$ the $\alpha$-Hölder rough path metric is defined as
[…]
Let $V^{\epsilon}_{t}=\dot{X}^{\epsilon}_{t}$ and $P^{\epsilon}_{t}=\epsilon^{2}V^{\epsilon}_{t}$ be velocity and momentum, respectively, then (1.1) is equivalent to
$\begin{cases}dX^{\epsilon}_{t}=V^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dV^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})V^{\epsilon}_{t}dt+\frac{1}{\epsilon^{2}}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon^{2}}dW_{t},\kern 5.0ptV^{\epsilon}_{0}=0,\end{cases}$ (2.1)
and
$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}P^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dP^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})P^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t},\kern 5.0ptP^{\epsilon}_{0}=0.\end{cases}$ (2.2)
Here we also introduce $Y^{\epsilon}_{t}:=\frac{1}{\epsilon}P^{\epsilon}_{t}$ [7, Page 7944], so that
$\begin{cases}dX^{\epsilon}_{t}=\frac{1}{\epsilon}Y^{\epsilon}_{t}dt,\kern 5.0ptX^{\epsilon}_{0}=0,\\ dY^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+\frac{1}{\epsilon}F(X^{\epsilon}_{t})dt+\frac{1}{\epsilon}dW_{t},\kern 5.0ptY^{\epsilon}_{0}=0.\end{cases}$ (2.3)
For each $0<\epsilon\leq 1,$ define the lift of $X^{\epsilon}$ by
$\mathbb{X}^{\epsilon}_{s,t}:=\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u},\quad(s,t)\in\Delta_{[0,T]},$
where the stochastic integral is Itô integral, and set $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon}).$ It is straightforward to check that $\mathbf{X}^{\epsilon}$ satisfies Chen relation. Moreover, by Kolmogorov criterion for rough paths [8, Theorem 3.1] and Proposition 3.3 below, $\mathbf{X}^{\epsilon}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
We impose the following assumptions:
($\mathbf{A1}$) There exists a constant $\lambda>0$, such that the smallest eigenvalue $\lambda_{1}$ of $\frac{M+M^{\top}}{2}$ satisfies that
$\lambda_{1}(x)\geq\lambda,\quad x\in\mathbb{R}^{d}.$
($\mathbf{A2}$) $M\in C^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$, and $M^{-1}\in C^{3}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$.
($\mathbf{A3}$) $M$, $M^{-1}$ and $\partial_{x_{j}}M^{-1}_{ik}$ are globally Lipschitz, and $M^{-1}$ is bounded, $1\leq i,j,k\leq d$.
($\mathbf{A4}$) $F$ is globally Lipschitz and bounded.
Remark 2.1.
From ($\mathbf{A1}$), one finds that
$y^{\top}M(x)y\geq\lambda|y|^{2},\quad(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d},$
which is essential in moment estimates. ($\mathbf{A2}$) is used to obtain regularity of the solution of Poisson equation in Section 4. Well-posedness of (1.2) and (2.3) are guaranteed by ($\mathbf{A3}$) and ($\mathbf{A4}$).
Now we state the main result.
Theorem 2.2.
Let $X$ be the solution of (1.2). Define
$\mathbb{X}_{s,t}:=\int_{s}^{t}X_{s,u}\otimes\circ dX_{u}+\frac{1}{2}\int_{s}^{t}J(X_{u})(M^{-1})^{\top}(X_{u})-M^{-1}(X_{u})J(X_{u})du,$ (2.4)
where $J$ is the solution of (1.3), and $\otimes\circ$ denotes the Stratonovich integral. Under assumptions $(\mathbf{A1})$$(\mathbf{A4})$, for each $\frac{1}{3}<\alpha<\frac{1}{2}$ and $1\leq p<\infty,$
$\lim\limits_{\epsilon\to 0}\mathbb{E}[\rho_{\alpha}(\mathbf{X}^{\epsilon},\mathbf{X})^{p}]=0,$
where $\mathbf{X}=(X,\mathbb{X}).$
Remark 2.3.
By Kolmogorov criterion for rough paths again, one verifies that $\mathbf{X}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
We end this section with a collection of frequently used notation.
Notation
$|\cdot|$: Euclidean norm of either a vector or a matrix.
$A^{\top}$: Transpose of a matrix or a vector $A$.
$\mathrm{id}$: Identity matrix.
$(x,y)$: Inner product of two vectors $x$ and $y$.
$Df,D^{2}f$: Gradient and Hessian of a function $f:\mathbb{R}^{d}\to\mathbb{R}$.
$C^{k}(U;V)$: The set consists of $k$-th continuously differentiable functions from $U$ to $V$, where $k\in\mathbb{N},$ and $U$, $V$ are Banach spaces.
$a\lesssim b$: There exists a constant $C>0$ such that $a\leq Cb.$ Throughout the paper, $C$ never depends on $\epsilon$. Additionally, in the proof of Proposition 3.3, it neither depends on $s$ nor $t$.
3 Moment Estimates
In this section, we give several moment estimate for system (2.3). We begin with bounding the supremum norm.
Proposition 3.1.
Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), For each $1\leq p<\infty,$
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty,\kern 5.0pt\sup\limits_{0<\epsilon\leq 1}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}|^{p}<\infty.$
Proof.
We firstly show the boundedness for $Y^{\epsilon}$. For each $p\geq 2,$ set $\phi(y)=|y|^{p},$ then
$D\phi(y)=p|y|^{p-2}y,\quad D^{2}\phi(y)=p(p-2)|y|^{p-4}yy^{\top}+p|y|^{p-2}\mathrm{id}.$
Apply Itô formula to $\phi(Y^{\epsilon}_{t})$ and use ($\mathbf{A1}$) and ($\mathbf{A4}$) to find that
$|Y^{\epsilon}_{t}|^{p}$ (3.1)
$=$ $\int_{0}^{t}[D\phi(Y^{\epsilon}_{s})]^{\top}dY^{\epsilon}_{s}+\frac{1}{2}\int_{0}^{t}\mathrm{Tr}\Big(\frac{1}{\epsilon^{2}}D^{2}\phi(Y^{\epsilon}_{s})\Big)ds$
$=$ $-\frac{p}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}(Y^{\epsilon}_{s})^{\top}M(X^{\epsilon}_{s})Y^{\epsilon}_{s}ds+\frac{p}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}(Y^{\epsilon}_{s})^{\top}F(X^{\epsilon}_{s})ds+H^{\epsilon}_{t}$
$+\frac{1}{2\epsilon^{2}}\int_{0}^{t}\mathrm{Tr}[p(p-2)|Y^{\epsilon}_{s}|^{p-4}Y^{\epsilon}_{s}(Y^{\epsilon}_{s})^{\top}+p|Y^{\epsilon}_{s}|^{p-2}\mathrm{id}]$
$\lesssim$ $-\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-1}ds+H^{\epsilon}_{t}+\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}ds,$
where $H^{\epsilon}$ is a martingale. Taking expectation and applying Hölder inequality,
$\mathbb{E}|Y^{\epsilon}_{t}|^{p}\lesssim-\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p-2}ds.$ (3.2)
First for $p=2$, by the comparison principle,
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{2}<\infty.$
By (3.2) and induction,
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty.$
for all even $p$, and hence for all $1\leq p<\infty.$
Next, we show the boundedness for $X^{\epsilon}$. Consider the following random ODE in $\mathbb{R}^{d\times d}$
$\dot{\phi}^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})\phi^{\epsilon}_{t},\kern 5.0pt\phi^{\epsilon}_{0}=\mathrm{id}.$
For each $r\in[0,T],$ set $z^{\epsilon,r}_{s}=\phi^{\epsilon}_{s}(\phi^{\epsilon}_{r})^{-1},\kern 5.0pts\in[r,T].$ It is showed in [27, Page 6] that, under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$),
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}dW_{r}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})drds.$ (3.3)
and that
$|z^{\epsilon,r}_{s}|\leq e^{-\epsilon^{-2}\lambda(s-r)},\quad 0\leq r\leq s\leq T.$ (3.4)
For notation simplicity, we always assume that $\lambda=1$. Then, from (3.4)
$\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|\leq\int_{r}^{t}|z^{\epsilon,r}_{s}|ds\leq\epsilon^{2}.$ (3.5)
Notice that by stochastic Fubini theorem, we have
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr.$ (3.6)
An application of Burkholder–Davis–Gundy inequality and (3.5) yield
$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}\Big|^{p}$ (3.7)
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{T}\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|^{2}dr\Big)^{p/2}$
$\lesssim$ $1.$
By ($\mathbf{A4}$) and (3.5), for each $t\in[0,T],$
$\frac{1}{\epsilon^{2p}}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}$ (3.8)
$\leq$ $\frac{1}{\epsilon^{2p}}\Big(\int_{0}^{t}\int_{r}^{t}|z^{\epsilon,r}_{s}|dsdr\Big)^{p}$
$\leq$ $T^{p}.$
Taking supremum and expectation,
$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}\lesssim 1.$ (3.9)
The result follows from combining (3.6), (3.7) and (3.9). ∎
We need the following elementary inequality, which is a special case of Minkowski integral inequality[10, Page 13]. For completeness, we give its proof.
Lemma 3.2.
Let $G$ be a random process on $[0,T],$ $a,b]\subset[0,T],$ $q\in\mathbb{N},$ then
$\Big[\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}\Big]^{1/q}\leq\int_{a}^{b}\Big(\mathbb{E}|G(r)|^{2q}\Big)^{1/q}dr.$
Proof.
Applying Hölder inequality,
$\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}$
$=$ $\mathbb{E}\int_{[a,b]^{q}}|G(r_{1})|^{2}\cdots|G(r_{q})|^{2}dr_{1}\cdots dr_{q}$
$=$ $\int_{[a,b]^{q}}\mathbb{E}(|G(r_{1})|^{2}\cdots|G(r_{q})|^{2})dr_{1}\cdots dr_{q}$
$\leq$ $\int_{[a,b]^{q}}(\mathbb{E}|G(r_{1})|^{2q})^{1/q}\cdots(\mathbb{E}|G(r_{q})|^{2q})^{1/q}dr_{1}\cdots dr_{q}$
$=$ $\Big(\int_{a}^{b}(\mathbb{E}|G(r)|^{2q})^{1/q}dr\Big)^{q}.$
The next proposition concerns about Hölder norm of the path $X^{\epsilon}$ and its lift.
Proposition 3.3.
Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), there is a constant $C$ independent of $\epsilon$, such that
(i) For each $1\leq p<\infty,$
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\leq C|t-s|^{p/2},\quad 0\leq s\leq t\leq T.$
(ii) For each $1\leq p<\infty,$
$\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u}\Big|^{p}\leq C|t-s|^{p},\quad 0\leq s\leq t\leq T.$
Proof.
In view of Jensen inequality, it is sufficient to prove the proposition for all even $p.$ We first prove (i). From (3.3) and stochastic Fubini theorem,
$X^{\epsilon}_{t}-X^{\epsilon}_{s}$ (3.10)
$=$ $\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu$
$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big]$
$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big].$
Now we deal with four terms above respectively. By Burkholder–Davis–Gundy inequality,
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$ (3.11)
From (3.4) and Hölder inequality
$\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}$
$\leq$ $\Big|\int_{s}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}$
$\leq$ $(t-s)\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du.$
Therefore,
$\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\lesssim(t-s)\int_{0}^{s}\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}dudr\lesssim\epsilon^{4}(t-s).$ (3.12)
Substituting the above inequality into (3.11),
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\epsilon^{2p}|t-s|^{p/2}.$ (3.13)
By Burkholder–Davis–Gundy inequality and (3.4),
$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$ (3.14)
$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}z^{\epsilon,r}_{u}du\Big|^{2}\Big)^{p/2}$
$\leq$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}dr\Big)^{p/2}$
$\leq$ $\Big(\int_{s}^{t}\epsilon^{4}dr\Big)^{p/2}$
$\leq$ $\epsilon^{2p}(t-s)^{p/2}.$
On the other hand, by ($\mathbf{A4}$), Hölder inequality and (3.12),
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.15)
$\lesssim$ $\mathbb{E}\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\mathbb{E}\Big\{\Big[\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}\Big\}$
$\lesssim$ $\mathbb{E}\Big\{\Big[\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big]^{p/2}\Big\}$
$\lesssim$ $\epsilon^{2p}|t-s|^{p/2}.$
Lastly,
$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.16)
$\lesssim$ $\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}dudr\Big|^{p}$
$\leq$ $\Big|\int_{s}^{t}\epsilon^{2}dr\Big|^{p}$
$\leq$ $\epsilon^{2p}(t-s)^{p}.$
Substituting (3.13)–(3.16) into (3.10),
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\lesssim|t-s|^{p/2}.$ (3.17)
We turn to prove (ii). By triangular inequality, it is sufficient to consider the $(i,j)$-entry $\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}$ of $\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u}.$ But from (3.3),
$dX^{\epsilon,j}_{t}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)dB^{k}_{r}dt+\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)F_{k}(X^{\epsilon}_{r})drdt,$
where $z^{\epsilon,r}_{t}(j,k)$ is the ($j,k$)-entry of $z^{\epsilon,r}_{t}$, so
$\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)dB^{k}_{r}du+\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)F_{k}(X^{\epsilon}_{r})drdu.$
Without loss of generality, we assume that the system is one-dimensional, that is, we write
$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}=\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu.$ (3.18)
By stochastic Fubini theorem,
$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}$ (3.19)
$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big]$
$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big]$
$=:$ $I+J.$
We firstly estimate $I$. Using Burkholder–Davis–Gundy inequality,
$\mathbb{E}|I|^{p}$ (3.20)
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$
Applying Lemma 3.2 to $G(r)=\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{0}^{s}\Big(\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.21)
Using Hölder inequality and taking $p$-th power,
$\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}\Big)^{p/2}.$ (3.22)
Applying Lemma 3.2 to $G(u)=|X^{\epsilon}_{s,u}|$ and $q=p/2,$ and then using (3.17),
$\Big\{\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big\}^{2/p}$
$\leq$ $\int_{s}^{t}(\mathbb{E}|X^{\epsilon}_{s,u}|^{p})^{2/p}du$
$\lesssim$ $\int_{s}^{t}|u-s|du$
$=$ $\frac{(t-s)^{2}}{2}.$
Taking ($p/2$)-th power,
$\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\lesssim|t-s|^{p}.$ (3.23)
On the other hand, from (3.4) we have
$\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\leq\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du\leq\frac{\epsilon^{2}}{2}e^{-2\epsilon^{-2}(s-r)}.$
Taking $(p/2)$-th power,
$\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\Big)^{p/2}\lesssim\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}.$ (3.24)
Combining (3.22), (3.23) and (3.24),
$\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}|t-s|^{p}.$
Taking ($2/p$)-th power and substituting into (3.21),
$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}$
$\lesssim$ $\int_{0}^{s}\epsilon^{2}e^{-2\epsilon^{-2}(s-r)}|t-s|^{2}dr$
$=$ $\epsilon^{2}(t-s)^{2}\int_{0}^{s}e^{-2\epsilon^{-2}(s-r)}dr$
$\leq$ $\epsilon^{4}(t-s)^{2}.$
Taking ($p/2$)-th power on the above inequality,
$\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\lesssim\epsilon^{2p}(t-s)^{p}.$ (3.25)
We deal with the other term in a similar fashion. Applying Lemma 3.2 to $G(r)=\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.26)
For fixed $r\in[s,t],$ using Lemma 3.2 to $G(u)=\sqrt{|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|},u\in[r,t]$ and $q=p$,
$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}$ (3.27)
$\leq$ $\Big(\mathbb{E}\Big[\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big]^{p}\Big)^{1/p}$
$\leq$ $\int_{r}^{t}\Big(\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p})^{1/p}du.$
By (3.4) and (3.17),
$\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}\mathbb{E}|X^{\epsilon}_{s,u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}|u-s|^{p/2}.$
Taking $(1/p)$-th power and substituting it into (3.27),
$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}\leq\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}|u-s|^{1/2}du\leq\epsilon^{2}|t-s|^{1/2}.$ (3.28)
Taking square and substituting it into (3.26),
$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\epsilon^{4}|t-s|dr=\epsilon^{4}|t-s|^{2}.$ (3.29)
Taking $(p/2)$-th power,
$\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\leq\epsilon^{2p}|t-s|^{p}.$ (3.30)
Combining (3.20), (3.25) and (3.30),
$\mathbb{E}|I|^{p}\lesssim|t-s|^{p}.$ (3.31)
It remains to deal with $J$. Writing
$\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.32)
$\lesssim$ $\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\Big[\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
$\lesssim$ $\Big[\int_{0}^{s}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
and
$\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.33)
$\lesssim$ $\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\Big[\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
$\lesssim$ $\Big[\int_{s}^{t}\Big(\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
so that we notice that the estimate of $J$ is identical to $I$ in view of (3.20), from which the proof is ended. ∎
An immediate consequence of Proposition 3.3 and Kolmogorov criterion for rough paths [8, Theorem 3.1] is the following corollary.
Corollary 3.4.
Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), for all $1\leq p<\infty$ and $\frac{1}{3}<\alpha<\frac{1}{2},$
[…]
Remark 3.5.
From Corollary 3.4, we see that $(X^{\epsilon})_{0<\epsilon\leq 1}$ is tight on $C([0,T];\mathbb{R}^{d}),$ which is assumed in [12]. But we do not use tightness throughout this paper.
4 Averaging Principle
In this section, we explore the averaging principle for system (2.3). To be precise, we establish the following proposition.
Proposition 4.1.
Let f be the form of either
$f(x,y)=x_{i}y_{k}y_{l}g(x),\quad x=(x_{1},...,x_{d})\in\mathbb{R}^{d},\kern 5.0pty=(y_{1},...,y_{d})\in\mathbb{R}^{d},$ (4.1)
or
$f(x,y)=y_{k}y_{l}g(x),\quad x=(x_{1},...,x_{d})\in\mathbb{R}^{d},\kern 5.0pty=(y_{1},...,y_{d})\in\mathbb{R}^{d},$ (4.2)
where $g\in C^{2}(\mathbb{R}^{d};\mathbb{R})$ is Lipschitz. Under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$), for each $1\leq p<\infty,$
$\lim\limits_{\epsilon\to 0}\mathbb{E}\Big|\int_{0}^{t}f(X^{\epsilon}_{s},Y^{\epsilon}_{s})-\bar{f}(X^{\epsilon}_{s})ds\Big|^{p}=0,$
where $\bar{f}(x):=\int f(x,y)\nu^{x}(dy)$ and $\nu^{x}$ is the unique invariant measure of the following system:
$dY^{x}_{t}=-M(x)Y^{x}_{t}dt+dW_{t}.$ (4.3)
Remark 4.2.
One sees that the non-linearity $F$ vanishes in (4.3). Intuitively speaking, the reason is that, the fast system in (2.3) with frozen slow variable
$dY^{\epsilon,x}_{t}=-\frac{1}{\epsilon^{2}}M(x)Y^{\epsilon,x}_{t}dt+\frac{1}{\epsilon}F(x)dt+\frac{1}{\epsilon}dW_{t},$
becomes
$d\tilde{Y}^{\epsilon,x}_{t}=-M(x)\tilde{Y}^{\epsilon,x}_{t}dt+\epsilon F(x)dt+d\tilde{W}_{t}$ (4.4)
after change-of-variable $\tilde{Y}^{\epsilon,x}_{t}:=Y^{\epsilon,x}_{\epsilon^{2}t},$ where $\tilde{W}_{t}:=\frac{1}{\epsilon}W_{\epsilon^{2}t}\overset{d}{=}W_{t},$ and the term $\epsilon F(x)$ in (4.4) disappears as $\epsilon$ goes to 0.
Remark 4.3.
It is known that [24, Proposition 3.5], under assumption ($\mathbf{A1}$), equation (4.3) admits a unique invariant measure which is Gaussian: $\nu^{x}=\mathcal{N}(0,J(x))$, where $J$ is the solution of (1.3). Consequently,
$\bar{f}(x)=x_{i}g(x)J_{kl}(x)$ (4.5)
if $f$ is of the form (4.1) and
$\bar{f}(x)=g(x)J_{kl}(x)$ (4.6)
if $f$ is of the form (4.2).
Remark 4.4.
It is known that [1, Page 179], under assumption ($\mathbf{A1}$), the Lyapunov equation
$M(x)A(x)+A(x)M^{\top}(x)=B(x)$
admits a unique solution
$A(x)=-\int_{0}^{\infty}e^{-M^{\top}(x)t}B(x)e^{-M(x)t}dt.$
This in particular implies that
$|A(x)|\leq\int_{0}^{\infty}|e^{-M^{\top}(x)t}|\cdot|B(x)|\cdot|e^{-M(x)t}|dt\leq\int_{0}^{\infty}e^{-2\lambda t}|B(x)|dt=\frac{1}{2\lambda}|B(x)|.$
Therefore, if $B$ is of polynomial growth, then so is $A$.
Proof of Proposition 4.1:
We only prove this proposition for $f$ of the form (4.1), since the other case is similar and easier. For each $x\in\mathbb{R}^{d}$, denote the generator of SDE (4.3) by $\mathcal{L}^{x},$ so that
$\mathcal{L}^{x}\phi(x,y)=(D_{y}\phi(x,y),-M(x)y)+\frac{1}{2}\mathrm{Tr}(D_{yy}\phi(x,y)),\quad\phi\in C^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d};\mathbb{R}).$ (4.7)
Consider Poisson equation
$\mathcal{L}^{x}\phi(x,y)=f(x,y)-\bar{f}(x).$ (4.8)
From (4.5) it is straightforward to check that
$\phi(x,y)=x_{i}g(x)(y^{\top}A(x)y-\mathrm{Tr}[A(x)J(x)]),$ (4.9)
where $J$ is the solution of (1.3) and $A$ is the solution of the following Lyapunov equation
$M^{\top}(x)A(x)+A(x)M(x)=\frac{1}{2}(e_{k}e_{l}^{\top}+e_{l}e_{k}^{\top}).$ (4.10)
Here, $(e_{n})_{1\leq n\leq d}$ is the orthonormal basis of $\mathbb{R}^{d}.$ As a consequence of ($\mathbf{A1}$)–($\mathbf{A3}$), $\phi\in C^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d};\mathbb{R}).$ By Itô formula and (4.7),
$d\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})$
$=$ $\Big(D_{x}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),\frac{1}{\epsilon}Y^{\epsilon}_{t}\Big)dt+\Big(D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}+\frac{1}{\epsilon}F(X^{\epsilon}_{t})\Big)dt$
$+\frac{1}{2}\mathrm{Tr}\Big[\frac{1}{\epsilon^{2}}D^{2}_{yy}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})\Big]dt+\Big(\frac{1}{\epsilon}D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),dW_{t}\Big)$
$=$ $\frac{1}{\epsilon}(D_{x}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),Y^{\epsilon}_{t})dt+\frac{1}{\epsilon}(D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),F(X^{\epsilon}_{t}))dt$
$+\frac{1}{\epsilon}(D_{y}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t}),dW_{t})+\frac{1}{\epsilon^{2}}\mathcal{L}^{X^{\epsilon}_{t}}\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})dt.$
Integrating on $[0,t]$, rearranging and using (4.8),
$\int_{0}^{t}f(X^{\epsilon}_{s},Y^{\epsilon}_{s})-\bar{f}(X^{\epsilon}_{s})ds$ (4.12)
$=$ $-\int_{0}^{t}\mathcal{L}^{X^{\epsilon}_{s}}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s})ds$
$=$ $-\epsilon^{2}[\phi(X^{\epsilon}_{t},Y^{\epsilon}_{t})-\phi(0,0)]+\epsilon\int_{0}^{t}(D_{x}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s}),Y^{\epsilon}_{s})ds$
$+\epsilon\int_{0}^{t}(D_{y}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s}),F(X^{\epsilon}_{s}))ds+\epsilon\int_{0}^{t}(D_{y}\phi(X^{\epsilon}_{s},Y^{\epsilon}_{s}),dW_{s}).$
With the aid of Proposition 3.1, the proof is finished by checking that $\phi$, $\phi_{x}$ and $\phi_{y}$ are of at most polynomial growth, that is, there exist constants $C>0$ and $q\geq 1$, such that for all $x,y\in\mathbb{R}^{d}$,
$\max\{|\phi(x,y)|,|\phi_{x}(x,y)|,|\phi_{y}(x,y)|\}\leq C(1+|x|^{q}+|y|^{q}).$
Differentiating on both sides of (4.9),
$\phi_{y}(x,y)=2x_{i}g(x)A(x)y,$ (4.13)
and
$\phi_{x_{j}}(x,y)$ (4.14)
$=$ $[\delta_{ij}g(x)+x_{i}\partial_{x_{j}}g(x)]\times[y^{\top}A(x)y-\mathrm{Tr}(A(x)J(x))]$
$+x_{i}g(x)\Big\{y^{\top}[\partial_{x_{j}}A(x)]y-\mathrm{Tr}[(\partial_{x_{j}}A(x))J(x)+A(x)\partial_{x_{j}}J(x)]\Big\}.$
Differentiating on both sides of (1.3) and (4.10), one finds that $\partial_{x_{j}}J(x)$ and $\partial_{x_{j}}A(x)$ satisfy Lyapunov equations
$M(x)\partial_{x_{j}}J(x)+\partial_{x_{j}}J(x)M^{\top}(x)=-([\partial_{x_{j}}M(x)]J(x)+J(x)[\partial_{x_{j}}M^{\top}(x)]).$ (4.15)
and
$M^{\top}(x)\partial_{x_{j}}A(x)+\partial_{x_{j}}A(x)M(x)=-([\partial_{x_{j}}M^{\top}(x)]A(x)+A(x)[\partial_{x_{j}}M(x)]),$ (4.16)
so the polynomial growth property is obtained by combining (4.9), (4.13), (4.14) and Remark 4.4.
5 Proof of the Main Theorem
With preparation above, we prove the main theorem now.
Proof of Theorem 2.2:
By Corollary 3.4 and [9, Theorem A.15], it suffices to prove the pointwise convergence, that is,
$X^{\epsilon}_{t}\to X_{t},\quad\text{in}\kern 5.0ptL^{p}(\Omega,\mathcal{F},P)$
and
$\mathbb{X}^{\epsilon}_{0,t}\to\mathbb{X}_{0,t},\quad\text{in}\kern 5.0ptL^{p}(\Omega,\mathcal{F},P)$
for each $0\leq t\leq T.$
Indeed, from (2.3),
$\epsilon dY^{\epsilon}_{t}=-\frac{1}{\epsilon}M(X^{\epsilon}_{t})Y^{\epsilon}_{t}dt+F(X^{\epsilon}_{t})dt+dW_{t}=-M(X^{\epsilon}_{t})dX^{\epsilon}_{t}+F(X^{\epsilon}_{t})dt+dW_{t}.$
Rearranging,
$dX^{\epsilon}_{t}=M^{-1}(X^{\epsilon}_{t})dW_{t}+M^{-1}(X^{\epsilon}_{t})F(X^{\epsilon}_{t})dt-\epsilon M^{-1}(X^{\epsilon}_{t})dY^{\epsilon}_{t},$ (5.1)
from which
$X^{\epsilon}_{t}$ $=$ $\Big[\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})dW_{s}+\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})ds\Big]-\epsilon\int_{0}^{t}M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}$ (5.2)
$=:$ $\mathbf{I}(\epsilon,t)-\mathbf{J}(\epsilon,t).$
Plainly,
$\mathbf{I}^{i}(\epsilon,t)=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds.$ (5.3)
By integration-by-part formula,
$\mathbf{J}^{i}(\epsilon,t)$ (5.4)
$=$ $\epsilon\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
$=$ $\epsilon\sum_{k}[M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}dM^{-1}_{ik}(X^{\epsilon}_{s})]$
$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\epsilon\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}$
$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds.$
Combining (5.2), (5.3) and (5.4),
$X^{\epsilon,i}_{t}$
$=$ $\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds$
$+\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds-\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}.$
From (1.2),
$X^{i}_{t}=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds+\sum_{k,l}\int_{0}^{t}J_{kl}(X_{s})\partial_{x_{l}}M^{-1}_{ik}(X_{s})ds,$
so
$X^{\epsilon,i}_{t}-X^{i}_{t}$ (5.5)
$=$ $\sum_{k}\int_{0}^{t}[M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})]dW^{k}_{s}+\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
$-\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\sum_{k,l}\int_{0}^{t}[Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})-J_{kl}(X^{\epsilon}_{s})\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})]ds$
$+\sum_{k,l}\int_{0}^{t}[J_{kl}(X^{\epsilon}_{s})\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})-J_{kl}(X_{s})\partial_{x_{l}}M^{-1}_{ik}(X_{s})]ds$
$=:$ $\mathbf{K}_{1}(\epsilon,t)+\mathbf{K}_{2}(\epsilon,t)+\mathbf{K}_{3}(\epsilon,t)+\mathbf{K}_{4}(\epsilon,t)+\mathbf{K}_{5}(\epsilon,t).$
By Burkholder–Davis–Gundy inequality, ($\mathbf{A3}$) and Hölder inequality
$\mathbb{E}|\mathbf{K}_{1}(\epsilon,t)|^{p}$ (5.6)
$\lesssim$ $\sum_{k}\mathbb{E}\Big|\int_{0}^{t}[M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})]dW^{k}_{s}\Big|^{p}$
$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})|^{2}ds\Big)^{p/2}$
$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{2}ds\Big)^{p/2}$
$\lesssim$ $d\mathbb{E}\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{p}ds$
$\lesssim$ $\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$
($\mathbf{A3}$) and ($\mathbf{A4}$) imply that
$|\mathbf{K}_{2}(\epsilon,t)|$
$\leq$ $\sum_{k}\int_{0}^{t}|M^{-1}_{ik}(X^{\epsilon}_{s})|\cdot|F_{k}(X^{\epsilon}_{s})-F_{k}(X_{s})|+|F_{k}(X_{s})|\cdot|M^{-1}_{ik}(X^{\epsilon}_{s})-M^{-1}_{ik}(X_{s})|ds$
$\lesssim$ $\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|ds.$
By Hölder inequality,
$\mathbb{E}|\mathbf{K}_{2}(\epsilon,t)|^{p}\lesssim\mathbb{E}\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{p}ds=\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$ (5.7)
By the same argument,
$\mathbb{E}|\mathbf{K}_{5}(\epsilon,t)|^{p}\lesssim\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds.$ (5.8)
Utilizing Proposition 3.1 and ($\mathbf{A3}$),
$\mathbb{E}|\mathbf{K}_{3}(\epsilon,t)|^{p}\lesssim\epsilon^{p}\sum_{k}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon,k}_{t}|^{p}\lesssim\epsilon^{p}.$ (5.9)
Lastly, from Proposition 4.1 and Remark 4.3, there exists a function $O(\epsilon,t)$ such that
$\mathbb{E}|\mathbf{K}_{4}(\epsilon,t)|^{p}\lesssim O(\epsilon,t),$ (5.10)
and that for each $t\in[0,T],$
$\lim\limits_{\epsilon\to 0}O(\epsilon,t)=0.$
Combining (5.5)–(5.10), we see that
$\mathbb{E}|X^{\epsilon,i}_{t}-X^{i}_{t}|^{p}\lesssim\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds+\epsilon^{p}+O(\epsilon,t).$
Summing up with respect to $i$,
$\mathbb{E}|X^{\epsilon}_{t}-X_{t}|^{p}\lesssim\int_{0}^{t}\mathbb{E}|X^{\epsilon}_{s}-X_{s}|^{p}ds+\epsilon^{p}+O(\epsilon,t).$
By Gronwall inequality,
$\lim\limits_{\epsilon\to 0}\mathbb{E}|X^{\epsilon}_{t}-X_{t}|^{p}=0.$ (5.11)
This shows the convergence of $X^{\epsilon}$. By Proposition 3.1 and dominated convergence theorem,
$\lim\limits_{\epsilon\to 0}\mathbb{E}\Big|\int_{0}^{t}\bar{f}(X^{\epsilon}_{s})-\bar{f}(X_{s})ds\Big|^{p}=0.$ (5.12)
In conjunction with Proposition 4.1,
$\lim\limits_{\epsilon\to 0}\mathbb{E}\Big|\int_{0}^{t}f(X^{\epsilon}_{s},Y^{\epsilon}_{s})-\bar{f}(X_{s})ds\Big|^{p}=0.$ (5.13)
Next, we turn to show the convergence of $\mathbb{X}^{\epsilon}.$ By (5.1),
$\int_{0}^{t}X^{\epsilon}_{s}\otimes dX^{\epsilon}_{s}$ (5.14)
$=$ $\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dW_{s}]+\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})]ds\Big)-\epsilon\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]$
$=:$ $\mathbf{A}(\epsilon,t)-\mathbf{B}(\epsilon,t).$
We explore the limit of (5.14) as $\epsilon\to 0$ componentwise. For $\mathbf{A}(\epsilon,t),$ by (5.11), Burkholder–Davis–Gundy inequality and dominated convergence theorem with the help of Proposition 3.1,
$\mathbf{A}^{ij}(\epsilon,t)$ (5.15)
$=$ $\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dW_{s}]\Big)_{ij}+\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})]ds\Big)_{ij}$
$=$ $\int_{0}^{t}X^{\epsilon,i}_{s}\times[M^{-1}(X^{\epsilon}_{s})dW_{s}]^{j}+\int_{0}^{t}X^{\epsilon,i}_{s}\times[M^{-1}(X^{\epsilon}_{s})F(X^{\epsilon}_{s})ds]^{j}$
$=$ $\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})F_{k}(X^{\epsilon}_{s})ds$
$\to$ $\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds,$
where the limit is in $L^{p}(\Omega,\mathcal{F},P).$ For $\mathbf{B}(\epsilon,t),$ integrating by parts,
$\mathbf{B}^{ij}(\epsilon,t)$ (5.16)
$=$ $\epsilon\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]\Big)_{ij}$
$=$ $\epsilon\int_{0}^{t}X^{\epsilon,i}_{s}[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]^{j}$
$=$ $\epsilon\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
$=$ $\epsilon\sum_{k}\Big(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]\Big)$
$=:$ $\mathbf{B}^{ij}_{1}(\epsilon,t)-\mathbf{B}^{ij}_{2}(\epsilon,t).$
By Proposition 3.1 and ($\mathbf{A3}$), $(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t})_{0<\epsilon\leq 1}$ is bounded in $L^{p}(\Omega,\mathcal{F},P)$ for each $t\in[0,T],$ so that
$\mathbf{B}_{1}^{ij}(\epsilon,t)\to 0\quad\text{in}\quad L^{p}(\Omega,\mathcal{F},P).$ (5.17)
On the other hand, notice that
$\mathbf{B}_{2}^{ij}(\epsilon,t)$ (5.18)
$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]$
$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,i}_{s}+\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}X^{\epsilon,i}_{s}\sum_{l}[\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}]$
$=$ $\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})ds+\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}X^{\epsilon,i}_{s}\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})ds$
$=:$ $\mathbf{B}_{21}^{ij}(\epsilon,t)+\mathbf{B}_{22}^{ij}(\epsilon,t).$
By (5.13),
$\mathbf{B}_{21}^{ij}(\epsilon,t)\to\sum_{k}\int_{0}^{t}M^{-1}_{jk}(X_{s})J_{ik}(X_{s})ds,$ (5.19)
and
$\mathbf{B}_{22}^{ij}(\epsilon,t)\to\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$ (5.20)
where the two limits above are both in $L^{p}(\Omega,\mathcal{F},P).$ Combining (5.15), (5.17), (5.19) and (5.20), we conclude that
$\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes dX^{\epsilon}_{s}\Big)_{ij}$ (5.21)
$\to$ $\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})F_{k}(X_{s})ds$
$+\sum_{k}\int_{0}^{t}M^{-1}_{jk}(X_{s})J_{ik}(X_{s})ds+\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$
in $L^{p}(\Omega,\mathcal{F},P).$
The proof is finished after verifying that the right hand side of (5.21) coincides with $\mathbb{X}^{ij}_{0,t},$ where $\mathbb{X}$ is defined via (2.4). Let us calculate $\mathbb{X}^{ij}_{0,t}.$ Writing (1.2) componentwise, we see that
$\int_{0}^{t}X^{i}_{s}dX^{j}_{s}$
$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds+\sum_{k}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s},$
and
$[X^{i},X^{j}]_{t}=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
However, multipling both sides of (1.3) by $M^{-1}$ on the left and $(M^{-1})^{\top}$ on the right and then writing it componentwise, we see that
$\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds=\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})J_{kj}(X_{s})+J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
Combining three equalities above,
$\int_{0}^{t}X^{i}_{s}\circ dX^{j}_{s}$
$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
$+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\frac{1}{2}\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X_{s})J_{kj}(X_{s})+J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds.$
Plainly,
$\frac{1}{2}\Big(\int_{0}^{t}J(X_{s})(M^{-1})^{\top}(X_{s})-M^{-1}(X_{s})J(X_{s})ds\Big)_{ij}$
$=$ $\frac{1}{2}\sum_{k}\int_{0}^{t}J_{ik}(X_{s})M^{-1}_{jk}(X_{s})-M^{-1}_{ik}(X_{s})J_{kj}(X_{s})ds.$
Therefore,
$\Big(\int_{0}^{t}X_{s}\otimes\circ dX_{s}\Big)_{ij}+\frac{1}{2}\Big(\int_{0}^{t}J(X_{s})(M^{-1})^{\top}(X_{s})-M^{-1}(X_{s})J(X_{s})ds\Big)_{ij}$ (5.22)
$=$ $\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{ik}(X_{s})F_{k}(X_{s})ds$
$+\sum_{k}\int_{0}^{t}X^{i}_{s}M^{-1}_{jk}(X_{s})dW^{k}_{s}+\sum_{k}\int_{0}^{t}J_{ik}(X_{s})M^{-1}_{jk}(X_{s})ds,$
which coincides with (5.21).